# Simplicial complexes with many facets are vertex decomposable

### v1

@@ 4. Consequences and Concluding Remarks
We end with some additional comments and examples connected to our main result. We first give an example to show that the bound in Theorem 1.1 is optimal.
@@
We construct, for any $22\leq c\leq n-$ 2, a pure simplicial complex $\Delta$ on $n$ vertices that has codimension $c$ and $\binom{n}{c}-2c$ facets, and that is not Cohen–Macaulay. Since any vertex decomposable complex is also Cohen–Macaulay, this shows that the bound in Theorem 1.1 is optimal.
We start with the complete $(n-c-1)$-skeleton of the simplex on $n$ vertices, and obtain $\Delta$ by removing the facets containing the following vertices: all of $\{1,\ldots,n-c-2\},$ exactly one of $\{n-c-1,n-c\}$, and exactly one of $\{n-c+1,\ldots,n\}$. Then we have removed exactly $2c$ facets, and ${\rm link}_{\Delta}(\{1,\ldots,n-c-2\})$ is disconnected, so not Cohen–Macaulay. Since every link in a Cohen–Macaulay complex is also Cohen–Macaulay, $\Delta$ is not Cohen–Macaulay.
@@
In the situation where $\dim(\Delta)=1$, so that $c=n-2$, we may view $\Delta$ as a graph on vertex set $[n]$ and our main result has a graph theoretic interpretation. Since $\Delta$ is assumed to be pure, we consider graphs with no isolated vertex. In this context, example 4.1 yields the union of a complete graph with an isolated edge. Indeed, from Lemma 2.3 we know that a pure 1-dimensional complex is vertex-decomposable if and only if it is connected. Hence in this case Theorem 1.1 is equivalent to the following statement in graph theory: an edge-cut of the complete graph $K_{n}$ into nontrivial parts has size at least $42c=2n-$ 4. We refer to [4] for definitions from graph theory.
@@
On the one hand, we can delete the facets $\{1,2\}$ and $\{1,3\}$, resulting in a complex that is easily seen not to be a matroid complex. On the other hand, if we start with $K_{4}$ and delete the facets $\{1,2\}$ and $\{3,4\}$ we are left with a 4-cycle, which is not shifted. Indeed the class of shifted 1-dimensional complexes coincide with the class of threshold graphs, which are known to be chordal.
We can also recover a result similar to [11, Theorem 2.1]. In particular, if we assume $\Delta$ has “even more” facets, then $\Delta$ is automatically pure, so it must be vertex decomposable.
@@
Let $\Delta$ be a simplicial complex on vertex set $[n]$ of codimension $c$. Suppose $\Delta$ has at least $\binom{n}{c}-c$ facets of dimension $\dim(\Delta)=n-c-1$. Then $\Delta$ is pure.
@@
If $\Delta$ has a facet $F$ of dimension smaller than $\dim(\Delta)$, then $F$ can be extended to a set $S$ containing $\dim(\Delta)$ vertices. Namely, if $G$ is a facet of dimension $\dim(\Delta)$, then $F\not\subset G$, giving $|G\backslash F|\geq\dim(\Delta)-\dim(F)+1$. We obtain the set $S$ by adding $\dim(\Delta)-\dim(F)$ vertices from $G\backslash F$ to $F$. Now, $S$ is contained in $c+1$ sets of size $n-c$, none of which are facets of $\Delta$ (as $S$ properly contains the facet $F$). This is a contradiction to the assumption that there are at least $\binom{n}{c}-c$ facets of dimension $\dim(\Delta)=n-c-1$.
We now derive a strengthening of [11, Theorem 2.1]:
@@
Let $\Delta$ be a simplicial complex on vertex set $[n]$ of codimension $c$. If $\Delta$ has at least $\binom{n}{c}-c$ facets of dimension $\dim(\Delta)=n-c-1$, then $\Delta$ is vertex decomposable.
@@
By Lemma 4.4, $\Delta$ is a pure simplicial complex of codimension $c$ with at least $\binom{n}{c}-c\geq\binom{n}{c}-2c+1$ facets. Now apply Theorem 1.1. ∎
Terai and Yoshida showed that Stanley-Reisner rings having large multiplicities are Cohen-Macaulay [11, Theorem 3.1] by proving that the dual ideal of the Stanley-Reisner ideal of $\Delta$ has a linear resolution [11, Theorem 3.3]. We can also recover this dual version as a direct consequence of our main result by employing Stanley-Reisner theory and the properties of Alexander duality, in particular, Eagon and Reiner’s classification of square-free monomial ideals with a linear resolution. The following corollary may be of interest for those researchers interested in monomial ideals. We assume that the reader is familiar with the Stanley-Reisner correspondence; for undefined terminology, see [8]. The following result originally appeared as [11, Theorem 3.3].
@@
Let $I$ be a square-free monomial ideal of $R=\mathbb{K}[x_{1},\ldots,x_{n}]$ such that every generator has degree $c$. If $I$ has at least $\binom{n}{c}-2c+1$ generators, then $I$ has a linear resolution.
@@
Given a square-free monomial $m=x_{i_{1}}\cdots x_{i_{c}}$ of degree $c$, set $F_{m}=[n]\setminus\{i_{1},\ldots,i_{c}\}$. So, if $I=\langle m_{1},\ldots,m_{t}\rangle$ is a square-free monomial ideal where each generator has degree $c$, then $\Delta=\langle F_{m_{1}},\ldots,F_{m_{c}}\rangle$ is a pure simplicial complex of codimension $c$ with at least $\binom{n}{c}-2c+1$ facets. By Theorem 1.1, $\Delta$ is a vertex decomposable simplicial complex, and consequently, $\Delta$ is Cohen-Macaulay.
@@
There are other ways to detect the vertex decomposability of a simplicial complex $\Delta$ in terms of combinatorial data. In [7] Lasoń provides a sufficient condition for vertex decomposability in terms of the $f$-vector of $\Delta$ (more specifically the number of facets and ridges). To describe these results, suppose $\Delta$ is a pure simplicial complex on vertex set $[n]$, with dimension $\dim(\Delta)=d$ and codimension $\operatorname{{codim}}(\Delta)=c=n-d-1$. Recall that the $f$-vector of $\Delta$ is $f(\Delta)=(f_{0},f_{1},\ldots,f_{d})$ where $f_{i}$ is the number of faces of $\Delta$ of dimension $i$. Theorem 1.1 implies that if $f_{d}\geq\binom{n}{c}-2c+1$, then $\Delta$ is vertex decomposable.
We say that a simplicial complex $\Delta$ is extremal if $\Delta$ is pure of dimension $t$ with $r$ facets such that it has the minimum number of faces of dimension $(t-1)$ among all simplicial complexes of dimension $t$ and $r$-faces.
As shown by [7, Corollary 1], an extremal simplicial complex can be determined from its $f$-vector. The main result from [7] is the following.
@@
[7, Theorem 5] An extremal simplicial complex is vertex decomposable.
@@
$\{1,2,3\},\leavevmode\nobreak\ \{1,4,5\},\leavevmode\nobreak\ \{1,2,6\},\leavevmode\nobreak\ \{1,4,6\},\leavevmode\nobreak\ \{2,4,5\},\leavevmode\nobreak\ \{3,4,5\},\leavevmode\nobreak\ \{3,5,6\}$
and any other 9 facets of dimension 2, for a total of 16 facets of dimension 2. In this case, the codimension is $c=3$, and $1516\geq\binom{6}{3}-2\cdot 3+1=$ 15, so, this simplicial complex is vertex decomposable by our Theorem 1.1. By our construction, this simplicial complex has $f$-vector $(6,15,16)$ (our initial choice of facets forces this simplicial complex to contain all the $1$-dimensional faces). However, this $f$-vector does not satisfy [7, Corollary 1]. Indeed, there is a pure simplicial complex with $f$-vector $(6,14,16)$, namely the simplicial complex which contains all the facets of dimension two except $\{1,5,6\},\leavevmode\nobreak\ \{2,5,6\},\leavevmode\nobreak\ \{3,5,6\},$ and $\{4,5,6\}$ (this simplicial complex does not contain the face $\{5,6\}$). We see that our simplicial complex is not extremal, and hence one can not deduce that it is vertex decomposable from [7].
@@ Acknowledgements
The authors thank Hailong Dao for his feedback and suggestions. They also thank Martin Milanič for answering questions about the graph theory literature. This project began at the workshop “Interactions Between Topological Combinatorics and Combinatorial Commutative Algebra” held at the Banff International Research Station (BIRS) in April 2023. We thank BIRS and the organizers for the opportunity to work together. Dochtermann is partly supported by Simons Foundation Grant #$\#964659$. Nair is partly supported by Dao’s Simons Foundation Grant FND0077558. Van Tuyl’s research is supported by NSERC Discovery Grant 2019-05412. Woodroofe’s research is supported in part by the Slovenian Research Agency (research program P1-0285 and research projects N1-0160, J1-2451, J1-3003 and J1-50000).

### v2

@@ 4. Consequences and Concluding Remarks
We end with some additional comments and examples related to our main result. We first give an example to show that the bound in Theorem 1.1 is optimal.
@@
For any $22\leq c\leq n-$ 2, we construct a pure simplicial complex $\Delta$ on $n$ vertices that has codimension $c$ and $\binom{n}{c}-2c$ facets, and that is not Cohen–Macaulay. Since any vertex decomposable complex is also Cohen–Macaulay, this shows that the bound in Theorem 1.1 is optimal.
We start with the complete $(n-c-1)$-skeleton of the simplex on $n$ vertices, and obtain $\Delta$ by removing the facets containing the following vertices: all of $\{1,\ldots,n-c-2\}$ (this set is $\{1\}$ if $n-c-2=1$ and is empty if $n-c-2=0$), exactly one of $\{n-c-1,n-c\}$, and exactly one of $\{n-c+1,\ldots,n\}$. Then we have removed exactly $2c$ facets, and ${\rm link}_{\Delta}(\{1,\ldots,n-c-2\})$ is $1-$dimensional and disconnected, meaning $\Delta$ is not Cohen–Macaulay, by Reisner’s Criterion. Since every vertex decomposable complex is Cohen-Macaulay, $\Delta$ is not vertex decomposable.
@@
In the situation where $\dim\Delta=1$, so that $c=n-2$, we may view $\Delta$ as a graph on vertex set $[n]$ and our main result has a graph theoretic interpretation. Since $\Delta$ is assumed to be pure, we consider graphs with no isolated vertex. In this context, Example 4.1 yields a complex $\Delta$ that is the union of a complete graph with an isolated edge, and the link being considered is ${\rm link}_{\Delta}(\emptyset)=\Delta$. Indeed, from Lemma 2.3 we know that a pure $1$-dimensional complex is vertex decomposable if and only if it is connected. Hence in this case Theorem 1.1 is equivalent to the following graph theoretic statement: an edge-cut of the complete graph $K_{n}$ into nontrivial parts has size at least $42c=2n-$ 4. We refer to [4] for definitions from graph theory.
@@
On the one hand, we can delete the facets $\{1,2\}$ and $\{1,3\}$, resulting in a complex that is easily seen not to be a matroid complex. On the other hand, if we start with $K_{4}$ and delete the facets $\{1,2\}$ and $\{3,4\}$ we are left with a 4-cycle, which is not shifted. Indeed the class of shifted $1$-dimensional complexes coincide with the class of threshold graphs [8], which are known to be chordal [10].
We can also recover a result similar to [14, Theorem 2.1]. In particular, if we assume $\Delta$ has “even more” facets, then $\Delta$ is automatically pure, so it must be vertex decomposable.
@@
Let $\Delta$ be a simplicial complex on vertex set $[n]$ of codimension $c$. Suppose $\Delta$ has at least $\binom{n}{c}-c$ facets of dimension $\dim\Delta=n-c-1$. Then $\Delta$ is pure.
@@
If $\Delta$ has a facet $F$ of dimension smaller than $\dim\Delta$, then $F$ can be extended to a set $S$ containing $\dim\Delta$ vertices. Namely, if $G$ is a facet of dimension $\dim\Delta$, then $F\not\subset G$, giving $|G\backslash F|\geq\dim\Delta-\dim F+1$. We obtain the set $S$ by adding $\dim\Delta-\dim F$ vertices from $G\backslash F$ to $F$. Now, $S$ is contained in $c+1$ sets of size $n-c$, none of which are facets of $\Delta$ (as $S$ properly contains the facet $F$). This is a contradiction to the assumption that there are at least $\binom{n}{c}-c$ facets of dimension $\dim\Delta=n-c-1$. ∎
We now derive a strengthening of [14, Theorem 2.1]:
@@
Let $\Delta$ be a simplicial complex on vertex set $[n]$ of codimension $c$. If $\Delta$ has at least $\binom{n}{c}-c$ facets of dimension $\dim\Delta=n-c-1$, then $\Delta$ is vertex decomposable.
@@
If $c=0$, then $\Delta$ is a simplex and the result follows immediately. So, suppose that $c\geq 1$. By Lemma 4.4, $\Delta$ is a pure simplicial complex of codimension $c$ with at least $\binom{n}{c}-c\geq\binom{n}{c}-2c+1$ facets. Now apply Theorem 1.1. ∎
Terai and Yoshida showed that Stanley-Reisner rings having large multiplicities are Cohen-Macaulay [14, Theorem 3.1], by proving that the dual ideal of the Stanley-Reisner ideal of $\Delta$ has a linear resolution [14, Theorem 3.3]. We can also recover this dual version as a direct consequence of our main result by employing Stanley-Reisner theory and properties of Alexander duality, and in particular Eagon and Reiner’s classification of square-free monomial ideals with a linear resolution. The following corollary may be of interest for those researchers working with monomial ideals. We assume that the reader is familiar with the Stanley-Reisner correspondence; for undefined terminology, see [11]. The following result originally appeared as [14, Theorem 3.3].
@@
Let $I$ be a square-free monomial ideal of $R=\mathbb{K}[x_{1},\ldots,x_{n}]$ with $\mathbb{K}$ an arbitary field, and suppose that every generator has degree $c$. If $I$ has at least $\binom{n}{c}-2c+1$ generators, then $I$ has a linear resolution.
@@
Given a square-free monomial $m=x_{i_{1}}\cdots x_{i_{c}}$ of degree $c$, set $F_{m}=[n]\setminus\{i_{1},\ldots,i_{c}\}$. So, if $I=\langle m_{1},\ldots,m_{t}\rangle$ is a square-free monomial ideal where each generator has degree $c$, then $\Delta=\langle F_{m_{1}},\ldots,F_{m_{t}}\rangle$ is a pure simplicial complex of codimension $c$ with at least $\binom{n}{c}-2c+1$ facets. By Theorem 1.1, $\Delta$ is a vertex decomposable simplicial complex, and consequently, $\Delta$ is Cohen-Macaulay.
@@
There are other ways to detect the vertex decomposability of a simplicial complex $\Delta$ in terms of combinatorial data. In [9] Lasoń provides a sufficient condition for vertex decomposability in terms of the $f$-vector of $\Delta$ (more specifically the number of facets and ridges). To describe these results, suppose $\Delta$ is a pure simplicial complex on vertex set $[n]$, with dimension $\dim\Delta=d$ and codimension $\operatorname{{codim}}\Delta=c=n-d-1$. Recall that the $f$-vector of $\Delta$ is $f(\Delta)=(f_{0},f_{1},\ldots,f_{d})$ where $f_{i}$ is the number of faces of $\Delta$ of dimension $i$. Theorem 1.1 implies that if $f_{d}\geq\binom{n}{c}-2c+1$, then $\Delta$ is vertex decomposable.
We say that a simplicial complex $\Delta$ is extremal if $\Delta$ is pure of dimension $t$ with $r$ facets, and it has the minimum number of faces of dimension $(t-1)$ among all simplicial complexes with $r$ faces of dimension $t$. As shown by [9, Corollary 1], an extremal simplicial complex can be determined from its $f$-vector. The main result from [9] is the following.
@@
[9, Theorem 5] An extremal simplicial complex is vertex decomposable.
@@
$\{1,2,3\},~{}\{1,4,5\},~{}\{1,2,6\},~{}\{1,4,6\},~{}\{2,4,5\},~{}\{3,4,5\},~{}\{3,5,6\}$
and any other 9 facets of dimension 2, for a total of 16 facets of dimension 2. In this case, the codimension is $c=3$, and $1516\geq\binom{6}{3}-2\cdot 3+1=$ 15, so, this simplicial complex is vertex decomposable by our Theorem 1.1. By our construction, this simplicial complex has $f$-vector $(6,15,16)$ (our initial choice of facets forces this simplicial complex to contain all the $1$-dimensional faces). However, this $f$-vector does not satisfy [9, Corollary 1]. Indeed, there exists a pure simplicial complex with $f$-vector $(6,14,16)$, namely the simplicial complex which contains all the facets of dimension two except $\{1,5,6\},~{}\{2,5,6\},~{}\{3,5,6\},$ and $\{4,5,6\}$ (this simplicial complex does not contain the face $\{5,6\}$). We see that our simplicial complex is not extremal, and hence one can not deduce that it is vertex decomposable from [9].
As a final consequence, we mention a connection to geometrically vertex decomposable ideals, as first defined by Klein and Rajchgot [7]. While we do not reproduce the formal definition here, a geometrically vertex decomposable ideal is a generalization of the properties of the square-free monomial ideals that corresponds to a vertex decomposable simplicial complex via the Stanley-Reisner correspondence. As shown in [7, Proposition 2.14], under a suitable lexicographical monomial order $<$, in some cases we can determine if an ideal $I$ is a geometrically vertex decomposable ideal from its initial ideal ${\rm in}_{<}(I)$. In particular, if ${\rm in}_{<}(I)$ is the Stanley-Reisner ideal of a vertex decomposable simplicial complex, then $I$ is geometrically vertex decomposable. By combining Klein and Rajchgot’s result with our Theorem 1.1, we get a new technique to check if an ideal is geometrically vertex decomposable. Precisely, if there is an appropriate monomial order $<$ such that the initial ideal ${\rm in}_{<}(I)$ is the Stanley-Reisner ideal of a simplicial complex with “enough” facets, as determined by Theorem 1.1, then $I$ is geometrically vertex decomposable.
@@ Acknowledgements
The authors thank the anonymous referees for their helpful comments and improvements. They are grateful to Hailong Dao for his feedback and suggestions, and also thank Martin Milanič for answering questions regarding the graph theory literature. This project began at the workshop “Interactions Between Topological Combinatorics and Combinatorial Commutative Algebra” held at the Banff International Research Station (BIRS) in April 2023. We thank BIRS and the organizers for the opportunity to work together. Dochtermann is partly supported by Simons Foundation Grant #$\#964659$. Nair is partly supported by Dao’s Simons Foundation Grant FND0077558. Van Tuyl’s research is supported by NSERC Discovery Grant 2019-05412. Woodroofe’s research is supported in part by the Slovenian Research Agency (research program P1-0285 and research projects N1-0160, J1-2451, J1-3003 and J1-50000).